# Conformalized Late Fusion Multi-View Learning

### v1

@@ 3 Method
We now introduce a procedure for uncertainty quantification for late-stage fusion. We discuss the naive approach in Section 3.1 and its extension along with the associated coverage guarantees in Section 3.2. We then discuss the downstream use of the resulting prediction regions in Section 3.3.
3.1 Naive Approach
We now consider the setting described in the background section, where $K$ predictors $f_{1},...,f_{K}$ are trained on the separate views with a late fusion algorithm $\mathcal{F}$, which we assume to be fixed throughout this presentation. Note that the validity of the results presented below is independent of the predictive quality of such $f_{i}$. The efficiency, however, is highly dependent on it. In the standard application of conformal prediction, a score function $s(\mathcal{F}(x),y)$ is defined. For instance, if we assume $\mathcal{Y}$ is some metric space with metric $d$, as in a regression task, a naive choice of score function would be
$s(x,y)=d(\mathcal{F}(f_{1}(x),...,f_{K}(x)),y).$ (2)
Denoting the $\lceil(N_{\mathcal{C}}+1)(1-\alpha)\rceil/N_{\mathcal{C}}$ quantile of the score distribution over $\mathcal{D}_{C}$ as $\widehat{q}(\alpha)$, $\mathcal{C}(x)=\{y:s(x,y)\leq\widehat{q}(\alpha)\}$ is then marginally calibrated. Ideally, however, prediction regions $\mathcal{C}(x)$ should have the quality that, if a particular predictor has less uncertainty in its predictions, as is frequently true of settings leveraging late-stage fusion, upon routing to that predictor, the corresponding size of the prediction region should be smaller than if it had been routed to a different predictor. While the framing of Equation 2 does, in principle, support this property, it ultimately relies on defining an uncertainty-aware late fusion algorithm $\mathcal{F}$. In its typical form, however, $\mathcal{F}$ simply takes point predictions $f_{1}(x),...,f_{K}(x)$ in as input, meaning any uncertainty-awareness would need to be baked in a priori through domain knowledge of the uncertainties of the predictors $f_{1},...,f_{K}$, which can seldom be specified precisely, in turn sacrificing the downstream predictive efficiency of $\mathcal{C}(x)$.
For this reason, we seek to produce a data-driven approach of obtaining the uncertainties of the predictors by instead fusing the conformalized predictors. In this case, we first define a score function $s_{i}$ for each predictor $f_{i}$. For any $x$, therefore, we now have a multivariate score $s(x,y):=(s_{1}(x,y),...,s_{K}(x,y))$. A naive approach would then involve leveraging standard conformal prediction over a pre-defined map $g:\mathbb{R}^{K}\rightarrow\mathbb{R}$, e.g., $g(s)=\sum_{k=1}^{K}s_{k}$. As in the case of conformalizing the fused predictor directly, using a fixed $g$ fails to adapt to any disparities in uncertainties present across predictors or requires intimate knowledge of such uncertainties. We instead wish to provide a data-adaptive pipeline to automatically produce such a $g$. Formally, denoting by $\mathcal{C}(g):\mathcal{X}\rightarrow\mathcal{Y}$ the conformalized predictor resulting from calibrating with the score fusion function $g$ in some function class $\mathcal{G}$, the goal can be formulated as seeking $g^{*}:=\operatorname*{arg\,min}_{g\in\mathcal{G}}\mathbb{E}_{X}\left[\mathcal{L}\left(\mathcal{C}\left(g\right)\left(X\right)\right)\right]$. There is, however, no approach to solving this abstract formulation, since even the mere evaluation of $\mathbb{E}_{X}[\mathcal{L}(\mathcal{C}(g)(X))]$ for a fixed $g$ is intractable.
3.2 MVCP: Score Quantile Envelope
Since finding $g^{*}$ exactly is intractable, we propose a data-adaptive approach that leverages the notion of quantile envelopes from Section 2.4 over the scores $s\in\mathbb{R}^{K}$. Importantly, for the remainder of this discussion, we assume the score functions are non-negative, i.e., $s_{k}:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}_{+}$, which is typically the case as the score serves as a generalization of the classical idea of a residual. Intuitively, the coverage guarantee in this multivariate extension requires us to define a pre-ordering $\lesssim$ over $\mathbb{R}^{K}$, where we only lose the antisymmetric axiom of a total ordering. Upon defining such a pre-ordering, the coverage guarantee follows in using the appropriate quantile under this modified sorting.
@@ 3.2 MVCP: Score Quantile Envelope
Crucially, however, the problem of choosing this pre-ordering closely parallels that of choosing $g^{*}$, where a poorly chosen pre-ordering will result in overly conservative quantile envelopes and, hence, prediction regions. For instance, using a lexicographical ordering $\lesssim_{\mathrm{Lex}}$ will result in axis-aligned hyper-rectangular quantile regions and similarly a norm-based ordering $\lesssim_{\mathrm{Norm}}$, defined as $s_{1}\lesssim_{\mathrm{Norm}}s_{2}\iff||s_{1}||_{2}\leq||s_{2}||_{2}$, in hyperspherical regions. Rather than explicitly prescribing a pre-ordering, $\lesssim$ can be implicitly defined by prescribing an indexed family of nested sets $\{\mathcal{A}_{t}\}_{t\in\mathbb{R}}$, such that $\mathcal{A}_{t_{1}}\subset\mathcal{A}_{t_{2}}$ for $t_{1}\leq t_{2}$. We then define $\lesssim$ by stating $s_{1}\lesssim s_{2}$ if $\forall t$, $s_{2}\in\mathcal{A}_{t}\implies s_{1}\in\mathcal{A}_{t}$.
@@
While not strictly necessary, we restrict ourselves to the case where $\mathcal{A}_{t}$ is a family of sets defined by a fixed “base contour” that scales outward with $t$. Under this restriction, performing a comparison of $s_{1},s_{2}\in\mathbb{R}^{K}$ in the aforementioned fashion, i.e. checking if $s_{1}\lesssim s_{2}$, amounts to checking if $t(s_{1})\leq t(s_{2})$, where $t(s)$ is the $t$ for which $\mathcal{A}_{t}$ “intersects” $s$. Notably, $t(s)$ is precisely the aforementioned data-driven score fusion function $g(s)$ of interest. Defining a data-adaptive $g(s)$, therefore, reduces to having a data-driven approach for defining the base contour of $\mathcal{A}_{t}$. We additionally restrict this base contour to be a convex set; if such a set were permitted to be nonconvex, computing $t(s):=\min\{t\in\mathbb{R}:s\in\mathcal{A}_{t}\}$ would potentially be computationally expensive, increasing the difficulty in defining the resulting prediction regions. As demonstrated in Section 3.3, many downstream applications require such efficient computation and, hence, benefit from this restriction to convex sets.
@@
To have tight quantile regions, we formally wish for the pre-ordering to have the property that the quantile region given by $\widetilde{\mathcal{Q}}:=\{s\mid s\lesssim\widehat{q}\}$, where $\widehat{q}$ is the $1-\alpha$ empirical quantile of $\mathcal{S}_{C}$ under $\lesssim$, should have minimal Lebesgue measure. Notably, by construction, such a $\widetilde{\mathcal{Q}}$ will capture $1-\alpha$ points of $\mathcal{S}_{C}$. The problem of discovering an optimal pre-ordering can, thus, be equivalently stated as seeking to define the base contour of $\mathcal{A}_{t}$ to match that of the tightest $1-\alpha$ convex cover of $\mathcal{S}_{C}$.
@@
This final reframing, therefore, naturally motivates selecting the base contour as the $1-\alpha$ quantile envelope of $\mathcal{S}_{C}$. Using $\mathcal{S}_{C}$ to define $\mathcal{A}_{t}$ and in turn $\lesssim$, however, sacrifices the exchangebility of its points with future test scores $s^{\prime}$, as the very nature of ordering would change in swapping $s^{\prime}$ with any $s\in\mathcal{S}_{C}$. The goal, then, follows as seeking to define the base contour as the $1-\alpha$ quantile envelope of $\mathcal{S}_{C}$ without directly using $\mathcal{S}_{C}$. For this reason, we partition $\mathcal{S}_{C}=\mathcal{S}_{C}^{(1)}\cup\mathcal{S}_{C}^{(2)}$, where we define $\lesssim$ using $\mathcal{S}_{C}^{(1)}$ and compute the final $\widehat{q}$ over $\mathcal{S}_{C}^{(2)}$. Such a split is predicated on the assumption that the $1-\alpha$ quantile envelope defined over $\mathcal{S}_{C}^{(1)}$ resembles that of $\mathcal{S}_{C}^{(2)}$, implying the $|\mathcal{S}_{C}^{(1)}|$ should be sufficiently large as to capture this structure accurately.
@@
We now focus attention on defining the quantile envelope over $\mathcal{S}_{C}^{(1)}$, starting by selecting the projection directions $\{u_{m}\}$. The restriction of $s$ to the positive orthant induces a natural modification of Equation 1, namely where $u\in\mathcal{S}_{+}^{K-1}:=\mathcal{S}^{K-1}\cap\mathbb{R}_{+}^{K}$ instead of $\mathcal{S}^{K-1}$. Exact solution of finding an evenly distributed set of points over hyperspheres in arbitrary $n$-dimensional spaces is a classically difficult problem [48]. If $K=2$, we can solve this exactly; for $K>2$, we generate directions stochastically such that $U\sim\mathrm{Unif}(\mathcal{S}_{+}^{K-1})$ by drawing $V_{1},...,V_{M}\sim\mathcal{N}(0,I^{K\times K})$ and defining $U_{i}:=V_{i}^{|\cdot|}/\sqrt{V_{1}^{2}+...+V_{M}^{2}}$, where $v^{|\cdot|}$ denotes the component-wise absolute values of $v$.
@@
We now wish to define the quantile thresholds $\{\widetilde{q}_{m}\}$ for the selected directions to optimally capture $1-\alpha$ of $\mathcal{S}_{C}^{(1)}$. Naively taking the $1-\alpha$ quantile per projection direction $u_{m}$ results in joint coverage by $\widetilde{\mathcal{Q}}:=\bigcap_{m=1}^{M}H(u_{m},\widetilde{q}_{m})$ of $\mathcal{S}^{(1)}_{\mathcal{C}}$ to be $<1-\alpha$. To combat this, a straightforward fix is to replace the $1-\alpha$ quantile per direction instead with its Bonferroni-corrected $1-\alpha/M$ quantile. While valid, this approach produces overly conservative prediction regions. We, therefore, instead tune a separate $\beta\in(\alpha/M,\alpha)$ parameter via binary search, finding the maximum $\beta^{*}$ such that using the $\beta^{*}$ quantile per direction provides the overall desired coverage, i.e. $|\bigcap_{m=1}^{M}H(u_{m},\widetilde{q}_{m}(1-\beta^{*}))\cap\mathcal{S}^{(1)}_{\mathcal{C}}|/N_{\mathcal{C}_{1}}\in(1-\alpha,1-\alpha+\epsilon)$ for some fixed, small $\epsilon>0$. With this choice of $\{(u_{m},\widetilde{q}_{m})\}$, we have a defined pre-ordering, whose coverage guarantees are formally stated below and proven in Appendix B.
@@
Suppose $\mathcal{D}_{\mathcal{C}}:=\{(X_{i},Y_{i})\}_{i=1}^{N_{\mathcal{C}}}$ and $(X^{\prime},Y^{\prime})$ are exchangeable. Assume further that $K$ maps $s_{k}:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ have been defined and a composite $s(X,Y):=(s_{1}(X,Y),...,s_{K}(X,Y))$ is defined. Further denote by $\mathcal{S_{C}}$ the evaluation of $s(X,Y)$ on $\mathcal{D_{C}}$, namely $\mathcal{S_{C}}:=\{s(X_{i},Y_{i})\}_{(X_{i},Y_{i})\in\mathcal{D}_{\mathcal{C}}}$. For some $\alpha\in(0,1)$, given a pre-order $\lesssim$ in $\mathbb{R}^{K}$ induced by a collection of nested sets $\{\mathcal{A}_{t}\}_{t\geq 0}$ define $\mathcal{Q}(\alpha)=\{s\in\mathbb{R}^{K}:s\lesssim s_{\lceil(N_{\mathcal{C}}+1)(1-\alpha)\rceil}\}$. Then, denoting $\mathcal{C}(X):=\{y:s(X,y)\in\mathcal{Q}(\alpha)\}$, $\mathcal{P}_{X,Y}(Y^{\prime}\in\mathcal{C}(X^{\prime}))\geq 1-\alpha$.
To compute $\widehat{q}$, for each $s\in\mathcal{S}^{(2)}_{\mathcal{C}}$, we find $t^{*}(s)$, defined to be $\min\{t\in\mathbb{R}:s\in\bigcap_{m=1}^{M}H(u_{m},t\widetilde{q}_{m})\}$. This can be efficiently computed as $t^{*}(s)=\max_{m=1,...,M}(u_{m}^{\top}s/\widetilde{q}_{m})$. Denoting the $\lceil(N_{\mathcal{C}_{2}}+1)(1-\alpha)\rceil$-th largest $t^{*}(s)$ as $\widehat{t}$, $\widehat{q}_{m}:=\widehat{t}\widetilde{q}_{m}$ and $\widehat{\mathcal{Q}}:=\bigcap_{m=1}^{M}H(u_{m},\widehat{q}_{m})$. Note that, if the tightest quantile envelope was already discovered over $\mathcal{S_{C}}^{(1)}$, this adjustment factor $\widehat{t}\approx 1$.
Importantly, while the aforementioned procedure will necessarily result in convex regions $\widehat{\mathcal{Q}}(\alpha)$, this does not mean the downstream prediction regions in $\mathcal{Y}$ will be convex, discussed more fully in Section 3.3. Such flexibility is necessary, as many studies have demonstrated the need for nonconvex prediction regions for downstream utility, such as [26, 49, 47]. However, it is unsurprising such flexibility exists, as even a single scalar score $s_{1}(x,y)$ can produce nonconvex prediction regions.
Another point to emphasis is that this procedure does not require a fusion algorithm to be defined: this is a key conceptual separation between MVCP and traditional conformal prediction. In the classic approach, a fused predictor is conformalized; here, instead, the conformalized predictors are fused. As a result, there is no notion of a point prediction for MVCP: it only returns prediction regions. This is fundamentally unlike the traditional case, where a point prediction $\widehat{y}:=\mathcal{F}(x)$ is “wrapped around” by a surrounding prediction region. We present the full algorithm in Algorithm 1 and an accompanying visual walkthrough of the procedure in Appendix A.
Algorithm 1 MVCP: $\textsc{UnifHypersphere}(K)$ is a subroutine that samples $\sim\mathrm{Unif}(\mathcal{S}^{K-1})$.

### v2

@@ 3 Method
We now introduce a procedure for conformal score aggregation for ensembles. We discuss the approach and its coverage guarantees in Section 3.1. We then discuss the downstream use of its prediction regions in Section 3.2.
3.1 Multivariate Score Quantile
We wish to consider the setting typical of conformal model aggregation, as discussed in Section 2.4, in which predictors $f_{1}(x),...,f_{K}(x)$ and corresponding scores $s_{1}(x,y),...,s_{K}(x,y)$ are defined. We further assume a similar premise as that taken by [40], in which the individual scores are stacked into a multivariate score $s(x,y):=(s_{1}(x,y),...,s_{K}(x,y))$. A naive approach would then involve leveraging standard conformal prediction over a pre-defined map $g:\mathbb{R}^{K}\rightarrow\mathbb{R}$, e.g., $g(s)=\sum_{k=1}^{K}s_{k}$. Similar to the naive conformalization of an ensembled predictor discussed in Section 2.4, using a fixed $g$ fails to adapt to any disparities in uncertainties present across predictors or requires intimate knowledge of such uncertainties. We instead wish to provide a data-adaptive pipeline to automatically produce such a $g$.
Importantly, we heretofore assume the score functions are non-negative, i.e., $s_{k}:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}_{+}$, which is typically the case as the score serves as a generalization of the residual. We highlight that many of the details of the method presented below are geometric in nature and are more easily understood with the supplement of diagrams. We have, thus, provided an accompanying visual walkthrough of the procedure in Appendix A to clarify its presentation.
3.1.1 Score Partial Ordering
Intuitively, our method seeks to directly generalize the approach of split conformal, by “ordering” the collection of multivariate calibration scores and taking the $1-\alpha$ score under such an ordering to be a threshold $\widehat{\mathcal{Q}}$ with which prediction regions are then implicitly defined. Formally, the multivariate “ordering” is established as a pre-ordering $\lesssim$ over $\mathbb{R}^{K}$; a pre-ordering differs from a total ordering in that it need not satisfy the antisymmetric axiom of a total ordering. Roughly speaking, an “acceptance region,” so called as it serves as the criterion used to ultimately decide which $y$ are accepted into the prediction region, is then defined as $\widehat{\mathcal{Q}}:=\{s\mid s\lesssim\widehat{q}\}$, where $\widehat{q}$ is the $1-\alpha$ empirical quantile of $\mathcal{S}_{C}$ under $\lesssim$. Such a $\widehat{\mathcal{Q}}$ naturally generalizes the standard scalar acceptance interval of $[0,\widehat{q}]$ in the case of non-negative score functions. We briefly highlight the distinction between acceptance regions and prediction regions: the former are subsets of the space of multivariate scores, i.e. $\subset\mathbb{R}^{K}$, that ultimately define the criteria for retaining particular $y$ values in the prediction region whereas the latter are the subsets of the output space that ultimately have coverage guarantees, i.e. $\subset\mathcal{Y}$. The two, however, are directly related; in particular, for a fixed score $s(x,y)$, a larger acceptance region will result in a more conservative prediction region.
@@ 3.1.1 Score Partial Ordering
Crucially, therefore, the problem of choosing this pre-ordering closely parallels that of choosing $g$, where a poorly chosen pre-ordering will result in overly large acceptance regions and, hence, conservative prediction regions. For instance, using a lexicographical ordering $\lesssim_{\mathrm{Lex}}$ will result in axis-aligned hyper-rectangular acceptance regions. As a result, rather than manually prescribing a pre-ordering, we define $\lesssim$ in a data-driven fashion by prescribing an indexed family of nested sets $\{\mathcal{A}_{t}\}_{t\in\mathbb{R}}$, such that $\mathcal{A}_{t_{1}}\subset\mathcal{A}_{t_{2}}$ for $t_{1}\leq t_{2}$ and stating $s_{1}\lesssim s_{2}$ if $\forall t$, $s_{2}\in\mathcal{A}_{t}\implies s_{1}\in\mathcal{A}_{t}$.
@@
For a family of sets $\{\mathcal{A}_{t}\}_{t\in\mathbb{R}}$, we take each $\mathcal{A}_{t}$ to be the region of the positive orthant $\mathbb{R}_{+}^{K}$ bounded by the coordinate axes and an “outer frontier” parameterized by $t$. The shape of this outer frontier remains fixed over the family and is merely scaled outward from the origin with $t$. Under this choice, performing a comparison of $s_{1},s_{2}\in\mathbb{R}^{K}$ in the aforementioned fashion, i.e. checking if $s_{1}\lesssim s_{2}$, amounts to checking if $t(s_{1})\leq t(s_{2})$, where $t(s)$ is the smallest $t$ for which the outer frontier of $\mathcal{A}_{t}$ intersects $s$. Notably, $t(s)$ is precisely the aforementioned data-driven score fusion function $g(s)$ of interest. Defining a data-adaptive $g(s)$, therefore, reduces to having a data-driven approach for defining the outer frontier of $\mathcal{A}_{t}$. We restrict this outer frontier to be such that $\mathcal{A}_{t}$ is a convex set; if $\mathcal{A}_{t}$ were permitted to be nonconvex, computing $t(s):=\min\{t\in\mathbb{R}:s\in\mathcal{A}_{t}\}$ would potentially be computationally expensive. The benefits of such convexity are highlighted, for example, in Section 3.2.
@@
To have tight acceptance regions, we formally wish for the pre-ordering to have the property that the acceptance region given by $\widetilde{\mathcal{Q}}$ has minimal Lebesgue measure and captures $1-\alpha$ points of $\mathcal{S}_{C}$. The problem of discovering an optimal pre-ordering can, thus, be equivalently stated as seeking to define the outer frontier of $\mathcal{A}_{t}$ to match that of the tightest $1-\alpha$ convex cover of $\mathcal{S}_{C}$.
@@
This final reframing, therefore, naturally motivates selecting the outer frontier to be the $1-\alpha$ quantile envelope of $\mathcal{S}_{C}$. Using $\mathcal{S}_{C}$ to define $\mathcal{A}_{t}$ and in turn $\lesssim$, however, sacrifices the exchangeability of its points with future test scores $s^{\prime}$, as the very nature of ordering would change in swapping $s^{\prime}$ with any $s\in\mathcal{S}_{C}$. The goal, then, follows as seeking to define the outer frontier as the $1-\alpha$ quantile envelope of $\mathcal{S}_{C}$ without directly using $\mathcal{S}_{C}$. For this reason, we partition $\mathcal{S}_{C}=\mathcal{S}_{C}^{(1)}\cup\mathcal{S}_{C}^{(2)}$, where we define $\lesssim$ using $\mathcal{S}_{C}^{(1)}$ and compute the final $\widehat{q}$ over $\mathcal{S}_{C}^{(2)}$. Such a split is predicated on the assumption that the $1-\alpha$ quantile envelope defined over $\mathcal{S}_{C}^{(1)}$ resembles that of $\mathcal{S}_{C}^{(2)}$, implying the $|\mathcal{S}_{C}^{(1)}|$ should be sufficiently large as to capture this structure accurately.
@@
We now focus attention on defining the quantile envelope over $\mathcal{S}_{C}^{(1)}$ using a technique paralleling that described in Section 2.2. In particular, we start by selecting the projection directions $\{u_{m}\}$ of Equation 1. The restriction of $s$ to the positive orthant induces a natural modification of Equation 1, namely where $u\in\mathcal{S}_{+}^{K-1}:=\mathcal{S}^{K-1}\cap\mathbb{R}_{+}^{K}$ instead of $\mathcal{S}^{K-1}$. To best approximate Equation 1, we wish for $\{u_{m}\}$ to be uniformly distributed over $\mathcal{S}_{+}^{K-1}$; however, exactly finding an evenly distributed set of points over hyperspheres in arbitrary $n$-dimensional spaces is a classically difficult problem [41] If $K=2$, we can solve this exactly; for $K>2$, we generate directions stochastically such that $U\sim\mathrm{Unif}(\mathcal{S}_{+}^{K-1})$ by drawing $V_{1},...,V_{M}\sim\mathcal{N}(0,I^{K\times K})$ and defining $U_{i}:=V_{i}^{|\cdot|}/\sqrt{V_{1}^{2}+...+V_{M}^{2}}$, where $v^{|\cdot|}$ denotes the component-wise absolute values of $v$.
@@
We now wish to define the quantile thresholds $\{\widetilde{q}_{m}\}$ for the selected directions to optimally capture $1-\alpha$ of $\mathcal{S}_{C}^{(1)}$. Naively taking the $1-\alpha$ quantile per projection direction $u_{m}$ results in joint coverage by $\widetilde{\mathcal{Q}}:=\bigcap_{m=1}^{M}H(u_{m},\widetilde{q}_{m})$ of $\mathcal{S}^{(1)}_{\mathcal{C}}$ to be $<1-\alpha$. A straightforward fix is to replace the $1-\alpha$ quantile per direction instead with its Bonferroni-corrected $1-\alpha/M$ quantile. While valid, this approach produces overly conservative prediction regions. We, therefore, instead tune a separate $\beta\in(\alpha/M,\alpha)$ parameter via binary search, finding the maximum $\beta^{*}$ such that using the $\beta^{*}$ quantile per direction provides the overall desired coverage, i.e. $|\bigcap_{m=1}^{M}H(u_{m},\widetilde{q}_{m}(1-\beta^{*}))\cap\mathcal{S}^{(1)}_{\mathcal{C}}|/N_{\mathcal{C}_{1}}\in(1-\alpha,1-\alpha+\epsilon)$ for some fixed, small $\epsilon>0$. With this choice of $\{(u_{m},\widetilde{q}_{m})\}$, we have a defined pre-ordering, whose coverage guarantees are formally stated below and proven in Appendix B.
@@
Suppose $\mathcal{D}_{\mathcal{C}}:=\{(X_{i},Y_{i})\}_{i=1}^{N_{\mathcal{C}}}$ and $(X^{\prime},Y^{\prime})$ are exchangeable. Assume further that $K$ maps $s_{k}:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ have been defined and a composite $s(X,Y):=(s_{1}(X,Y),...,s_{K}(X,Y))$ is defined. Further denote by $\mathcal{S_{C}}$ the evaluation of $s(X,Y)$ on $\mathcal{D_{C}}$, namely $\mathcal{S_{C}}:=\{s(X_{i},Y_{i})\mid(X_{i},Y_{i})\in\mathcal{D}_{\mathcal{C}}\}$. For some $\alpha\in(0,1)$, given a pre-order $\lesssim$ in $\mathbb{R}^{K}$ induced by a collection of nested sets $\{\mathcal{A}_{t}\}_{t\geq 0}$ define $\mathcal{Q}(\alpha)=\{s\in\mathbb{R}^{K}:s\lesssim s_{\lceil(N_{\mathcal{C}}+1)(1-\alpha)\rceil}\}$. Then, denoting $\mathcal{C}(x):=\{y:s(x,y)\in\mathcal{Q}(\alpha)\}$, $\mathcal{P}_{X^{\prime},Y^{\prime}}(Y^{\prime}\in\mathcal{C}(X^{\prime}))\geq 1-\alpha$.
3.1.2 Score Quantile Threshold
To then compute $\widehat{q}$, we find $t^{*}(s)$ for each $s\in\mathcal{S}^{(2)}_{\mathcal{C}}$, defined to be $\min\{t\in\mathbb{R}:s\in\bigcap_{m=1}^{M}H(u_{m},t\widetilde{q}_{m})\}$. This can be efficiently computed as $t^{*}(s)=\max_{m=1,...,M}(u_{m}^{\top}s/\widetilde{q}_{m})$. Denoting the $\lceil(N_{\mathcal{C}_{2}}+1)(1-\alpha)\rceil$-th largest $t^{*}(s)$ as $\widehat{t}$, $\widehat{q}_{m}:=\widehat{t}\widetilde{q}_{m}$ and $\widehat{\mathcal{Q}}:=\bigcap_{m=1}^{M}H(u_{m},\widehat{q}_{m})$. If the tightest quantile envelope was already discovered over $\mathcal{S_{C}}^{(1)}$, this adjustment factor $\widehat{t}\approx 1$.
Importantly, while the aforementioned procedure will necessarily result in convex regions $\widehat{\mathcal{Q}}$, this does not mean the downstream prediction regions in $\mathcal{Y}$ will be convex, as discussed in Section 3.2. Such flexibility is necessary, as many studies have demonstrated the need for nonconvex prediction regions for downstream utility, such as [36, 42, 43]. However, it is unsurprising such flexibility exists, as even a single scalar score $s_{1}(x,y)$ can produce nonconvex prediction regions. We present the full algorithm in Algorithm 1.
Algorithm 1 CSA: $\textsc{UnifHypersphere}(K)$ is an assumed subroutine that samples $\sim\mathrm{Unif}(\mathcal{S}^{K-1})$.